Date: 02-01-26
Time: 0653hrs
Location: 1500 Nepperhan Ave x Barney St
Description: Fire throughout a 1.5-story dwelling with water problems.
0656hrs- Homeowner rpts fire from an outlet
E310, 309, 308, 312; L75, 72, R-1, Squad 11, B2
0659hrs- Additional calls for smoke in the area. B2 req FAST
L71 FAST
0659hrs- E310 has smoke showing from a distance
0700hrs- E310 on location with smoke from the attic
0702hrs- B-2 on location with smoke from the roof transmitting 10-29 (Working Fire)
0706hrs- B-2 rpts 2-story dwelling, no exposure problems. Transmitting 10-30 (All Hands)
0712hrs- Relocating E314, L70 to Station 12
0715hrs- B-2 transmitting 2nd Alarm, companies rpt water problems.
2nd Alarm: E306, 314, L70
0716hrs- B-2 rpts (3) LS, fire through the roof.
0722hrs- E314 to hit hydrant at Hearst X Nepperhan
0723hrs- E313, L73 to Station 12.
0729hrs- B-2 rpts removing all members and going to exterior operations
0750hrs- (2) master streams in operation, req. water department to increase pressure.
0757hrs- B-2 rpts shutting down TL75 and sending companies interior.
0802hrs- CMD rpts making progress
0818hrs- CMD rpts all co's being removed from structure due to roof collapse.
Time: 0653hrs
Location: 1500 Nepperhan Ave x Barney St
Description: Fire throughout a 1.5-story dwelling with water problems.
0656hrs- Homeowner rpts fire from an outlet
E310, 309, 308, 312; L75, 72, R-1, Squad 11, B2
0659hrs- Additional calls for smoke in the area. B2 req FAST
L71 FAST
0659hrs- E310 has smoke showing from a distance
0700hrs- E310 on location with smoke from the attic
0702hrs- B-2 on location with smoke from the roof transmitting 10-29 (Working Fire)
0706hrs- B-2 rpts 2-story dwelling, no exposure problems. Transmitting 10-30 (All Hands)
0712hrs- Relocating E314, L70 to Station 12
0715hrs- B-2 transmitting 2nd Alarm, companies rpt water problems.
2nd Alarm: E306, 314, L70
0716hrs- B-2 rpts (3) LS, fire through the roof.
0722hrs- E314 to hit hydrant at Hearst X Nepperhan
0723hrs- E313, L73 to Station 12.
0729hrs- B-2 rpts removing all members and going to exterior operations
0750hrs- (2) master streams in operation, req. water department to increase pressure.
0757hrs- B-2 rpts shutting down TL75 and sending companies interior.
0802hrs- CMD rpts making progress
0818hrs- CMD rpts all co's being removed from structure due to roof collapse.